OpenStudy (anonymous):

3(5z-7)-2(9z-11)=4(8z-13)-17

OpenStudy (saifoo.khan):

solve

OpenStudy (anonymous):

z=2

OpenStudy (saifoo.khan):

\[\large 15z - 21 - 18z + 22 = 32z - 52 - 17\]

OpenStudy (anonymous):

15z - 21 - 18z + 22 = 32z - 52 - 17 -3z + 1 = 32z - 69 70 = 35z z = 2

OpenStudy (anonymous):

type fast saifoo

OpenStudy (saifoo.khan):

-3z + 1 = 32z - 69 35z = 70 z = 70/35 \[\Large z = 2\]

OpenStudy (saifoo.khan):

ty!